OpenStudy (anonymous):

HELP ME (1/a+1)(1/b+1)(1/c+1)=1 and P=a^2+b^2+c^2

OpenStudy (perl):

what are the directions for the problem

OpenStudy (rational):

You want to find the value of a^2 + b^2 + c^2 ?

OpenStudy (perl):

@freckles interesting the question is a month old

OpenStudy (freckles):

but I see @Hackerpham5 here like he did bump and he is here so maybe he will still say something

OpenStudy (rational):

he gave me medal

OpenStudy (freckles):

maybe your guess was right then

OpenStudy (perl):

maybe the directions say "find P =a^2 + b^2 + c^2"

OpenStudy (freckles):

http://openstudy.com/users/hackerpham5#/updates/54e2ab86e4b0db7d6b0f5288 look there looks like there is a little more here?

OpenStudy (freckles):

I don't know what all that other stuff means though .

OpenStudy (freckles):

might be a small language barrier

OpenStudy (freckles):

find the min

OpenStudy (freckles):

for a,b,c of (0;1) and (1/a+1)(1/b+1)(1/c+1)=1 find the min P=a^2+b^2+c^2 is what that translates to

OpenStudy (freckles):

We can try to do lagrange multipliers maybe

OpenStudy (rational):